OpenStudy (anonymous):

the sum of the digits of two-digit number is 15. if the digist are reversed the new number is 27 less than the original number. Find the original number

OpenStudy (yuki):

interesting problem

OpenStudy (yuki):

let's say that the 10th digit is a and the units to be b

OpenStudy (yuki):

then a + b = 15

OpenStudy (yuki):

our two digit number can be represented as 10a + b

OpenStudy (yuki):

so if you reverse the order it becomes 10b + a

OpenStudy (yuki):

the new one is 27 less than the original number, so your other equation will be 10b+a = 10a+b -27

OpenStudy (yuki):

now that you have two equations with two unknowns, you can solve for the system of equations.

OpenStudy (anonymous):

please help me with the answer :(

OpenStudy (anonymous):

the two eqs are a+b=10 10b+a-10a-b=-27 9a-9b=-27

OpenStudy (radar):

which is it a+b=15 or a+b=10?

OpenStudy (anonymous):

ah...a+b=15

OpenStudy (anonymous):

Did you guys find the number?

OpenStudy (anonymous):

a+b=15 9a-9b=-27 are the two eqs solving the two eqs... a=6 n b=9

OpenStudy (radar):

96 , 9+6=15 69=96-27

OpenStudy (anonymous):

So, the original number is 96.

OpenStudy (radar):

Looks like a good Bingo lol

OpenStudy (anonymous):

but as yuki took the places of digits,,10a+b gives 69:S

OpenStudy (anonymous):

i guess second eq should be 10b+a+27=10a+b ???

OpenStudy (radar):

I gave yuki a medal anyway

OpenStudy (radar):

You're getting one too

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

n u know i made mistake in solving a qs too :P

OpenStudy (radar):

It is the spirit of the effort that counts.

OpenStudy (anonymous):

right...but it surly confuses the one who posts the question

OpenStudy (anonymous):

thank you guys u are life saaaavvveeeerrrssss